# NestQuant: Nested Lattice Quantization for Matrix Products and LLMs

### v1

@@ Appendix B Gosset oracle
2:$d\leftarrow\text{floor}(x)$
3:$x_{2}\leftarrow d+0.5$
4:$g\leftarrow(x>x_{2})$
5:$s\leftarrow 2\cdot g-1$
6:$x_{1}\leftarrow d+g$
7:$dist\leftarrow(x-x_{2})\cdot s$
8:$\Delta\leftarrow\sum_{i}dist_{i}$
@@
Appendix C Crude estimate of runtime of quantized matrix
We give an estimate of how the NestQuant runtime affects the overall efficiency of the model. We note that the runtime of encoding and decoding one 8-vector with one scaled codebook is dominated by running the oracle for the Gosset lattice, described in Algorithm 5. By counting the arithmetic operations and comparisons equally, we get that one run of the oracle takes approximately 15 operations per vector entry. We should also add no more than $3$ arithmetic operations in Algorithms 1 and 2, and multiplication by the generator matrix $G$, or its inverse $G^{-1}$. Given the structure of the matrices, we can spend less operations on multiplying them by vectors using prefix sum technique for $G^{-1}$ and ignoring zeros for $G$. So, the multiplication takes around $2$ operations per vector entry. In total, we do 20 FLOPs per dequantizing one matrix entry, and $20\cdot k$ FLOPs for quantization, since we need to run the algorithm for $k$ values of $\beta$. We note that in weight quantization the encoding time is not important since it’s done offline, and in KV cache quantization encoding happens one time per token, while decoding happens with any additional query to the KV cache.
$G^{-1}=\begin{pmatrix}0.5&0.5&0.5&0.5&0.5&0.5&0.5&-\tfrac{7}{2}\\ 0&1&1&1&1&1&1&-6\\ 0&0&1&1&1&1&1&-5\\ 0&0&0&1&1&1&1&-4\\ 0&0&0&0&1&1&1&-3\\ 0&0&0&0&0&1&1&-2\\ 0&0&0&0&0&0&1&-1\\ 0&0&0&0&0&0&0&2\end{pmatrix}\;\;\;G=\begin{pmatrix}2&-1&0&0&0&0&0&0.5\\ 0&1&-1&0&0&0&0&0.5\\ 0&0&1&-1&0&0&0&0.5\\ 0&0&0&1&-1&0&0&0.5\\ 0&0&0&0&1&-1&0&0.5\\ 0&0&0&0&0&1&-1&0.5\\ 0&0&0&0&0&0&1&0.5\\ 0&0&0&0&0&0&0&0.5\end{pmatrix}$ (4)
Recall that in generation phase the LLM runtime is memory-bound, i.e. getting the weights from memory takes significantly more time than computing the matrix-vector product. Let’s compute the total running time of loading matrix of size $m\times n$ from memory, which can be decomposed into the following running times:
@@ Appendix C Crude estimate of runtime of quantized matrix
    Loading the matrix from memory $mnRT_{\mathrm{load}}/8$ seconds, where $R$ is the rate (bits/entry) and $T_{\mathrm{load}}$ is the time to load a byte (in seconds)
@@
    Decoding the matrix $N_{\mathrm{decode}}\cdot mn\cdot T_{\mathrm{flop}}$ where $N_{\mathrm{decode}}$ is the number of operations per entry and $T_{\mathrm{flop}}$ is the time to apply a floating point operation to a byte.
In total $T_{\mathrm{total}}=mn(RT_{\mathrm{load}}/8+T_{\mathrm{flop}}N_{\mathrm{decode}})$ compared to $mn\cdot 2T_{\mathrm{load}}$ (assume 16-bit baseline), so the speedup is $\frac{2T_{\mathrm{load}}}{RT_{\mathrm{load}}/8+T_{\mathrm{flop}}N_{\mathrm{decode}}}$.
Defining the arithmetic intensity $\alpha=\frac{T_{\mathrm{load}}}{T_{\mathrm{flop}}}$, we get a theoretical speedup of $\frac{2\alpha}{R\alpha/8+N_{\mathrm{decode}}}$. In our implementation of quantization $N_{\mathrm{decode}}=20$ and $\alpha\approx 30$, For $R=4$ we should expect a theoretical speedup of $\frac{60}{35}\approx 1.71$. For $R=3$ we get a theoretical speedup of $\frac{62}{33.25}\approx 1.92$.
In summary, we expect NestQuant in to achieve speedup in two following ways:
  • By speeding up the retrieval of weights and KV cache entries
  • By quantizing activations for large models, where weights are stored on different GPUs, and there are bandwidth limitations.
Appendix D Dynamic programming for optimal $\beta$
@@
Given a set of 8-vectors to quantize, we can find the set of $\beta$ that minimizes reconstruction error using a dynamic programming algorithm, which is described in Appendix D.
@@
    First-$\beta$: Use the smallest $\beta$, which does not result in an overflow error.
@@
    Opt-$\beta$: Try all the values of $\beta$, and choose the one that has the smallest reconstruction MSE.
@@
Opt-$\beta$ 0.0878 0.0795 0.0708 0.0669 0.0646
First-$\beta$ 0.0878 0.0798 0.0712 0.0676 0.0656
Table 4: Mean RMSE for reconstructed iid standard Gaussian 8-vectors, $q=16$, $k$ betas are uniform on $[0,10]$.
Even though Opt-$\beta$ should provide smaller error, the definition of First-$\beta$ will be useful for us. We can note that the difference between error for Opt-$\beta$ and First-$\beta$ is not very siginificant (Table 4). Moreover, First-$\beta$ can be used to determine the optimal set of $\beta_{i}$ to use.
Let we have $n$ samples $v_{1},v_{2},\ldots,v_{n}$ from the distribution of vectors we are quantizing, and a large set of betas $B$, containing $\beta_{1}<\beta_{2}<\ldots<\beta_{m}$, from which we want to take the optimal subset of size $k$ which minimizes the loss under First-$\beta$ strategy. For each vector $v_{i}$ and beta $\beta_{j}$ we compute $mse_{ij}$ — the MSE if we use $\beta_{j}$ to quantize $v_{i}$ and $overload_{ij}$ — whether an overload error occurs in this scenario.
We solve the optimization problem with dynamic programming. Let’s define $dp_{ij}$ be the mimimum sum of MSE we can get if we have to quantize all the vectors which do not yield an overload error for $\beta_{i}$, using $\beta_{i}$ and $j-1$ smaller betas and First-$\beta$ strategy. If $i$ is large enough so that no vector has an overflow error on $\beta_{i}$, $dp_{ik}$ has the answer to the problem. To compute the value of $dp_{ij}$, we can iterate over $s$ — the index of second largest beta in the set (the largest being $\beta_{i}$). Then, the recalculation works in the following way:
@@
where $\text{where }cond_{p}$ $=overload_{ps}\wedge\neg overload_{pi}$
@@
2:$dp_{i,j}=\infty$ for $i$ in $0\ldots m$, $j$ in $0\ldots k$
3:$from_{i,j}=\text{null}$ for $i$ in $0\ldots m$, $j$ in $0\ldots k$
4:$dp_{0,0}=0$
@@
Appendix E Llama experiment details
We choose the train split of the Wikitext2 [31] dataset as a calibration dataset for computing $H$, and evaluate the model on the validation split, computing the perplexity metric. For step 2 in the algorithm (Section 4.5), we select $\hat{\beta}=[3.5,4.5,6.0,14.5,25.0]/q$, because it is the $\beta$ we get when optimizing them for weight quantization without consideration of LDLQ. The overall universe of betas contains values from $1$ to $40$ with spacing ranging from $0.25$ to $2$. For running DP on activations, keys, and values, we run the model on a batch of $6$ full-length sequences, which is sufficient for this low-dimensional hyperparameter.
When choosing maximum beta for given distribution, we add a margin of $\frac{3.0}{q}$ for weights and $\frac{4.0}{q}$ to the maximum beta needed to have 00 overload errors on known data to account for potential overload errors in unknown data. While small number of overload error does not affect perplxity significantly, we still aim to minimize their probability.
@@ Appendix E Llama experiment details
Appendix F Ablation studies
@@
While Hadamard matrices from Sylvester construction are commonly used in other works (QuIP#, Quarot), there are multiple ways to construct a fast rotation for the case when dimension is not a power of $2$ (such as the down projection in MLP of Llama-3). We tested three possible options for rotation on $q=14$, $k=4$, W + KV + A quantization.

### v3

@@ Appendix C Gosset oracle
[…]
@@
Appendix D NestQuantM algorithm
Due to the high complexity of $\operatorname*{\arg\!\min}$ and $\operatorname*{\arg\!\max}$ operations in algorithm 5 for the fast hardware implementation, we propose a different, more simple version of NestQuant decoding. Instead of using $\operatorname*{\arg\!\max}$ and $\operatorname*{\arg\!\min}$ on lines 10 and 15, we always assign $pos$ to be equal to 111 (thus, indicating that we always "flip" the rounding direction of the first coordinate to fix the parity). Note that this change is only applied in the decoding stage, while during encoding we use the full version of the algorithm, which keeps the granular error the same. Let’s denote our modified Gosset oracle as $f\colon\mathbb{R}^{8}\to\mathbb{R}^{8}$.
Lemma D.1.
For a vector $x\in\mathbb{R}^{8}$ and a vector $v\in E_{8}$, $f(x+v)=f(x)+v$.
Proof.
Recall that $D_{8}\subset\mathbb{Z}^{8}$ contains all the integer vector with even sum of coordinates. The modified Gosset oracle $f$ uses modified $D_{8}$ oracle $g$, and at input $x$ constructs two candidate points $c_{1}(x)\in D_{8}$ and $c_{2}(x)\in D_{8}+\frac{1}{2}$. Then, the algorithm chooses the closest point among these candidates to $x$. Note that $c_{1}(x)=g(x)$ and $c_{2}(x)=g\left(x-\frac{1}{2}\right)+\frac{1}{2}$.
We will prove that for $u\in D_{8}$, $g(x+u)=g(x)+u$ for any $x\in\mathbb{R}^{8}$. Now, let’s show the original lemma. Note that since we are choosing the closest candidate point, if the condition on candidate sets $\{c_{1}(x+v),c_{2}(x+v)\}=\{c_{1}(x),c_{2}(x)\}+v$ holds, then $f(x+v)=f(x)+v$. Now, consider two cases:
@@ Appendix D NestQuantM algorithm
    $v\in D_{8}$. Then:
    $c_{1}(x+v)$ $=g(x+v)=g(x)+v=c_{1}(x)+v$
    $c_{2}(x+v)$ $=g\left(x+v-\frac{1}{2}\right)+\frac{1}{2}=g\left(x-\frac{1}{2}\right)+\frac{1}{2}+v=c_{2}(x)+v$
@@
    $v\in D_{8}+\frac{1}{2}$. Then, we say that $v=u-\frac{1}{2}=w+\frac{1}{2}$ for $u,v\in D_{8}$.
    $c_{1}(x+v)$ $=g(x+v)=g\left(x+u-\frac{1}{2}\right)=g\left(x-\frac{1}{2}\right)+u=g\left(x-\frac{1}{2}\right)+\frac{1}{2}+v=c_{2}(x)+v$
    $c_{2}(x+v)$ $=g\left(x-\frac{1}{2}+v\right)+\frac{1}{2}=g(x+w)+\frac{1}{2}=g(x)+w+\frac{1}{2}=g(x)+v=c_{1}(x)+v$
Thus, in both cases the condition on candidate sets holds, and we get $f(x+v)=f(x)+v$.
Now we show that if $u\in D_{8}$, $g(x+u)=g(x)+u$. Note that when evaluating $g(x+u)$, we will get the same rounding directions as in $g(x)$, since $u$ is an integer vector. Since $u$ has even sum of coordinates, the parity will also be the same. Then, our decision to flip the rounding of the first coordinate will also match as well. Therefore, the vector between original and rounded coordinates will be the same:
$g(x)-x=g(x+u)-x-u\Rightarrow g(x+u)=g(x)+u$
Let $v$ be the vector we obtain after rounding to $E_{8}$, $c$ be the lattice coordinates of $v$, and $G$ be the generating matrix of the lattice. The compressed representation is $c\;\mathrm{mod}\ q$, which corresponds to a point $v^{\prime}=G(c\;\mathrm{mod}\ q)$. The reconstructed point $\hat{v}$ is defined to be $v^{\prime}-qf(v^{\prime}/q)$ by the decoding algorithm. Note that $v-v^{\prime}\in qE_{8}$. Let’s assume that $f(v/q)=0$. Then:
$\hat{v}=v^{\prime}-qf\left(\frac{v^{\prime}}{q}\right)=v^{\prime}-qf\left(\frac{v}{q}+\frac{v^{\prime}-v}{q}\right)=v^{\prime}-q\left(f\left(\frac{v}{q}\right)+\frac{v^{\prime}-v}{q}\right)=v^{\prime}-q\cdot\frac{v^{\prime}-v}{q}=v$
We have used Lemma D.1 in the third equality and our assumption in the fourth equality. Given this fact, we conclude that when using NestQuantM for decoding, the composition of encoding and decoding functions (for a fixed $\beta$) is similar to the original NestQuant, except the shaping region has changed to the set of points $v\in E_{8}$ such that $f(v/q)=0$. Since $f$ is close to the original Gosset oracle, we expect this region to still capture Gaussian probability density well. In case of overload errors, we are still able to choose a larger value of beta due to multi-beta strategy.
Appendix E CUDA Kernel Implementation
The decoding algorithm 2 consists of a basis change $p\leftarrow Gc$ and a subsequent computation of the coset of $\frac{p}{q}$,
$\frac{p}{q}-Q_{\Lambda}\left(\frac{p}{q}\right).$
We leverage the asymmetry between encoding and decoding choosing $G$ which is fast to decode. In particular we use
$G=\begin{pmatrix}1&0&0&0&0&0&0&0\\ 1&0&2&0&0&0&0&0\\ 1&0&0&0&2&0&0&0\\ 1&0&0&0&0&0&2&0\\ 1&4&2&2&2&2&2&2\\ 1&0&0&2&0&0&0&0\\ 1&0&0&0&0&2&0&0\\ 1&0&0&0&0&0&0&2\end{pmatrix}$
$G$ is not a basis for $E_{8}$ but for $2E_{8}$. We use $G$ because we want to work in integer-lattice as half-integers cannot be represented with integers. Moreover, we represent 8-vector $\vec{x}$ as two 32 bit integers each representing the two parts of the vector $x_{0},x_{1},x_{2},x_{3}$ and $x_{4},x_{5},x_{6},x_{7}$.
For reference, the CUDA implementation of the implementation of $p\leftarrow Gc$.
1__device__ void G_q_fast(
2 const uint32_t enc, uint32_t* out0, uint32_t* out1
3)
4{
5 uint32_t even = enc & 0x0F0F0F0F;
6 uint32_t odd = (enc & 0xF0F0F0F0) >> 4;
7
8 uint32_t two_even = even << 1;
9 uint32_t two_odd = odd << 1;
10 uint32_t x0 = even & 0xFF;
11 uint32_t concatenator = (1 << 24) | (1 << 16) | (1 << 8) | 1;
12 uint32_t x0_concat = x0 * concatenator;
13
14 uint32_t temp_even = __vadd4(two_even & 0xFFFFFF00, x0_concat);
15 uint32_t temp_odd = __vadd4(two_odd, x0_concat);
16
17 uint32_t s1 = __dp4a(two_even, (uint32_t)0x01010100, (uint32_t)0);
18 uint32_t s2 = __dp4a(two_odd, (uint32_t)0x01010101, s1);
19
20 *out0 = temp_even;
21 *out1 = temp_odd + s2;
22}
The computation is done by decomposing $G$ into a sum of different matrices.
$G$ $=\begin{pmatrix}1&0&0&0&0&0&0&0\\ 1&0&2&0&0&0&0&0\\ 1&0&0&0&2&0&0&0\\ 1&0&0&0&0&0&2&0\\ 1&4&2&2&2&2&2&2\\ 1&0&0&2&0&0&0&0\\ 1&0&0&0&0&2&0&0\\ 1&0&0&0&0&0&0&2\end{pmatrix}=\begin{pmatrix}1&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&0\\ 1&0&0&0&0&0&0&0\\ \end{pmatrix}+\begin{pmatrix}0&0&0&0&0&0&0&0\\ 0&0&2&0&0&0&0&0\\ 0&0&0&0&2&0&0&0\\ 0&0&0&0&0&0&2&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\end{pmatrix}$
$+\begin{pmatrix}0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&2&0&0&0&0&0&0\\ 0&0&0&2&0&0&0&0\\ 0&0&0&0&0&2&0&0\\ 0&0&0&0&0&0&0&2\\ \end{pmatrix}+\begin{pmatrix}0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&2&2&2&2&2&2&2\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0\end{pmatrix}$
The next step is to compute $\frac{p}{q}$ and the coset of $\frac{p}{q}$ we merge both operations into one function.
1__device__ void decode_nestquant(
2 uint32_t *enc, int *B_local_decode, const int N = 8
3)
4{
5 uint32_t inp0, inp1;
6 unsigned int dist[2];
7
8 const uint32_t MASK_FRACPART = 0x1F1F1F1F;
9 const uint32_t MASK_INTPART = 0xE0E0E0E0;
10 const int HALF_CONCAT = 0x10101010;
11 const int TWO = 0x40;
12 const int ONE = 0x20;
13 const int HALF = 0x10;
14
15 G_q_fast(enc[0], &inp0, &inp1);
16
17 int32_t fracPart0 = inp0 & MASK_FRACPART;
18 int32_t fracPart1 = inp1 & MASK_FRACPART;
19
20 int32_t integerPart0 = inp0 & MASK_INTPART;
21 int32_t integerPart1 = inp1 & MASK_INTPART;
22
23 int g0 = (fracPart0 & HALF_CONCAT) << 1;
24 int g1 = (fracPart1 & HALF_CONCAT) << 1;
25
26 int32_t change = ((g0 & ONE) - HALF);
27 int32_t sum1 = get_sum(integerPart0, integerPart1);
28 int32_t two_parity_1 = (sum1 & ONE) >> 4;
29 int32_t fracPart0_1 = __vsub4(fracPart0, (int32_t)(two_parity_1 * change) & 0xFF);
30
31 int32_t sum2 = get_sum(g0, g1) + sum1;
32 int32_t two_parity_2 = (sum2 & ONE) >> 4;
33 int32_t fracPart0_2 = __vadd4(fracPart0, (int32_t)(two_parity_2 * change) & 0xFF);
34
35 dist[0] = __vabsdiffs4(fracPart0, HALF_CONCAT);
36 dist[1] = __vabsdiffs4(fracPart1, HALF_CONCAT);
37 int Delta = get_sum(dist[0], dist[1]);
38
39 int32_t dist00 = (dist[0] & 0xFF);
40 Delta -= (two_parity_1 + two_parity_2) * dist00;
41 Delta += two_parity_1 << 4;
42
43 if (Delta <= TWO) {
44 B_local_decode[0] = __vsub4(fracPart0_1, HALF_CONCAT);
45 B_local_decode[1] = __vsub4(fracPart1, HALF_CONCAT);
46 } else {
47 B_local_decode[0] = __vsub4(fracPart0_2, g0);
48 B_local_decode[1] = __vsub4(fracPart1, g1);
49 }
50}
Remember that for $q=16$, we would need to divide $p$ by 161616. We used a basis for $2E_{8}$, and not $E_{8}$ so every element is doubled. This is why in our code $\frac{1}{2}$ is represented by 16, 111 is represented by 323232, and 222 is represented by 646464. In order to avoid using the round function explicitly (which is not parallelized in hardware) we compute the integer and fractional part of the vector using masks (as can be seen in lines 17-21).
We compute $g$ in two parts $g_{0}$ and $g_{1}$ (see lines 23 and 24). This is done by checking that the fractional part is greater or equal to $\frac{1}{2}$ (equivalently in our scaled basis, that the fractional part has the fourth LSB set to one). We leverage the fact that
$x-(\lfloor x\rfloor+0.5)=\{x\}-0.5$
(lines 44 and 45) and
$x-\mathrm{round}(x)=\lfloor x\rfloor+\{x\}-(\lfloor x\rfloor+g)=\{x\}-g$
(lines 47 and 48). To compute the bitflip, we compute $2g_{1}-1$ (corresponding to the first element in the vector, see line 26 for reference). We add (subtract) it from fractional part based on the parity check we compute in lines 28 and 32 respectively.
To decode betas in our kernel, we encode the betas as indices to a predefined dictionary (of size 4). Thus, each beta can be represented using 2 bits.
1int beta1 = (beta_packed >> shift) & 0x3;
2int beta2 = (beta_packed >> (shift + 2)) & 0x3;
3int decoded_beta1 = beta_dict[beta1];
4int decoded_beta2 = beta_dict[beta2];
Encoding and decoding kernels share the same logical structure so that the complex mapping defined for beta is shared (this allows fast contiguous reads of multiple betas at once).
E.1 Runtime comparison of GEMV
Table 4: Runtime comparison of GEMV kernels on an $\times 8192$ 8192 matrix using an NVIDIA A100 GPU.
Method Time (µs)
Baseline (16 bits) 97
NestQuantM (4.25 bits) 60
QuIP# (2 bits) 38
QuIP# (4 bits) $\sim$75
int4 uniform 31
The QuIP# computation involves invoking two calls to QuIP# (2 bits), so we extrapolate the running time based on QuIP# (2 bits).
Appendix F Dynamic programming for optimal $\beta$
@@
Given a set of 8-vectors to quantize, we can find the set of $\beta$ that minimizes reconstruction error using a dynamic programming algorithm, which is described in Appendix F.
@@
[…]
@@
[…]
@@
[…]
Table 5: Mean RMSE for reconstructed iid standard Gaussian 8-vectors, $q=16$, $k$ betas are uniform on [0,10][0,10][ 0 , 10 ].
[…]
@@
$\text{where }cond_{p}$ $=overload_{ps}\wedge\neg overload_{pi}$
@@
[…]
@@
Appendix G Llama experiment details
We choose the train split of the Wikitext2 [35] dataset as a calibration dataset for computing $H$, and evaluate the model on the validation split, computing the perplexity metric. For step 2 in the algorithm (Section 4.6), we select $\hat{\beta}=[3.5,4.5,6.0,14.5,25.0]/q$, because it is the $\beta$ we get when optimizing them for weight quantization without consideration of LDLQ. The overall universe of betas contains values from 111 to 404040 with spacing ranging from 0.250.250.25 to 222. For running DP on activations, keys, and values, we run the model on a batch of 666 full-length sequences, which is sufficient for this low-dimensional hyperparameter.
When choosing maximum beta for given distribution, we add a margin of $\frac{3.0}{q}$ for weights and $\frac{4.0}{q}$ to the maximum beta needed to have 0 overload errors on known data to account for potential overload errors in unknown data. While small number of overload error does not affect perplxity significantly, we still aim to minimize their probability.
@@ Appendix G Llama experiment details
Appendix H Ablation studies
@@
While Hadamard matrices from Sylvester construction are commonly used in other works (QuIP#, Quarot), there are multiple ways to construct a fast rotation for the case when dimension is not a power of 222 (such as the down projection in MLP of Llama-3). We tested three possible options for rotation on $q=14$, $k=4$, W + KV + A quantization.
@@
H.1 The choice of $k$
Appendix I Results for Llama3.2-1B
q Bits Bits (no zstd) W W + KV W + KV + A
14 3.99 4.06 10.061 10.529 11.197
12 3.76 3.837 10.178 10.862 11.910
10 3.50 3.57 10.377 11.552 14.191
8 3.18 3.25 10.850 13.309 18.710
Appendix J Results for 3-bit model quantization
We publish the results for 3-bit quantization of weights and activations on small models (Llama-3-8B and Llama-2-7B). We use $q=7$ and $k=4$, which results in 2.98 bits per entry.
Bits (W-A-KV) Method Llama-2-7B Llama-3-8B
16-16-16 Floating point 5.47 6.14
4-4-16 NestQuant 5.64 6.56
3-3-16 NestQuant 8.25 6.33